OpenStudy (anonymous):

Luanne made a photocopy of a drawing of a rectangle. The original drawing was 4.5 inches wide and 6 inches long. The copy was 6.75 inches wide and 9 inches long. What was the scale factor for luannnes copy? Show your work: A.1.5 B.2 C. 2.25 D. 3

jagr2713 (jagr2713):

the scale factor is: 6.75/4.5=9/6=1.5

OpenStudy (anonymous):

A. I think